solve by the elimination method
6x-9y = 25.5
7y-3x= -14
The solution is (type an ordered pair, simplify your answer)

Answer by checkley77(12844)

6x-9y = 25.5
7y-3x= -14 multiply by 2 & add.
14y-6x=-28
-9y+6x=25.5
------------------
5y=-2.5
y=-2.5/5
y=-.5 ans.
6x-9(-.5)=25.5
6x+4.5=25.5
6x=25.5-4.5
6x=21
x=21/6 ans.
Proof:
7(-.5)-3*21/6=-14
-3.5-63/6=-14
-3.5-10.5=-14
-14=-14